OpenStudy (anonymous):

Reid and Maria both play soccer. This season, Reid scored 4 less than twice the number of goals that Maria scored. The difference in the number of goals they scored was 6. How many goals did each of them score?

OpenStudy (anonymous):

m = maria's goals r = reid's goals r = 2m - 4 m - r = 6 @ivettef365 .....take a peek at this please.....what am I doing wrong...I am getting a negative number

OpenStudy (anonymous):

did I set it up wrong ?

OpenStudy (anonymous):

@amistre64 ....I am feeling real stupid right now....can you help

OpenStudy (ivettef365):

set up is good then r = m-6

OpenStudy (ivettef365):

I got r = 16 and m = 10

OpenStudy (ivettef365):

what is wrong is that r -m = 6

OpenStudy (anonymous):

I did that because there was a difference of 6 in their scores.

OpenStudy (ivettef365):

yes, but Reid goes first

OpenStudy (anonymous):

if r = 2m - 4 and r = m - 6 then.. m - 6 = 2m - 4.....???

OpenStudy (ivettef365):

r = 2m - 4 r -m = 6 so r = m+6 m+6 = 2m - 4 4 + 6 = 2m -m so m = 10

OpenStudy (anonymous):

gotcha......I feel stupid

OpenStudy (ivettef365):

I got confuse at first :)

OpenStudy (anonymous):

thanks for coming and straightening everything out :)

OpenStudy (ivettef365):

not a problem :)

OpenStudy (amistre64):

the difference is an absolute value .... the difference between 5 and 2 is: 5 - 2 = 3 2 - 5 = -3 but since we are not concerned with direction ... just magnitude, then |5-2| = |2-5| = 3